A rocket burns fuel at a rate of 251 kg/s and
exhausts the gas at a relative speed of 7 km/s.
Find the thrust of the rocket.
Answer in units of MN.
Don't round answer.
1
Expert's answer
2019-12-05T10:19:06-0500

The thrust force of the rocket


"F=\frac{\Delta m}{\Delta t}v"

"=251\:\rm kg/s\times 7000\: m/s=1.757\times 10^6 \: N"